5) Solve the quadratic equation using the quadratic formula: 2x² - 5 = 4

6. Solve using the quadratic formula:
x² – 3x – 3 = -5
thank you so much

Answer by checkley71(8405)

2x^2-5=4
2x^2-5-4=0
2x^2-9=0 using the quadratic equation x=(-b+-sqrt[b^2-4ac])/2a we get
x=(0+-sqrt[0^2-4*2*-9])/2*2
x=(+-sqrt[72])/4
x=8.485/4
x=2.12 solution
x=-8.485/4
x=-2.12 solution
----------------------------------------------------
x^2-3x-3=-5
x^2-3x-3+5=0
x^2-3x+2=0
x=(3+-sqrt[-3^2-4&1*2])/2*1
x=(3+-sqrt9-8])/2
x=(3+-sqrt1)/2
x=(3+-1)/2
x=(3+1)/2
x=4/2
x=2 solution
x=(3-1)/2
x=2/2
x=1 solution